# Adjoint motives of modular forms and the Tamagawa number conjecture

### v1

@@ 6.5. Exceptional primes
where $V_{p}^{\prime}$ denotes the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to II mod $p^{c_{p}/2}$, and $\Theta(\varepsilon)$ is defined in §3.2 of [C-D-T]. Our hypotheses ensure that $\varepsilon/\varepsilon\circ\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ has conductor $(p{\mathcal{O}}_{F})^{c_{p}/2}$. Let $g_{p}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{c_{p}/2}}}\end{array}\right)$}}$. We recall the following properties of $\Theta(\varepsilon)$:
@@
where ${\mathcal{V}}^{\Sigma}_{p,\Sigma_{0}}={\mathcal{V}}_{p}^{0}$ if $p\in\Sigma_{0}-\Sigma$ and ${\mathcal{V}}_{p}^{1}$ (restricted to $U_{p}$) otherwise. Here the tensor product is over all primes $p$, but the action is trivial for all but finitely many pp. Note that if $\Sigma_{0}\subset\Sigma$, then $U^{\Sigma}_{\Sigma_{0}}=U^{\Sigma}$ and $\sigma^{\Sigma}_{\Sigma_{0}}=\sigma(N,\psi)$.
@@
It follows from the lemma and the definitions that $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$ is injective and commutes with the action of the Hecke operators $T_{p}$ for all primes pp. In particular, for $p\in S$, the operator $T_{p}$ can be written as a polynomial in $T_{q}$ for $q\neq p$, and the same arguments as in Proposition 5.6 show that $\mathbb{T}$} acts on ${\mathcal{M}}(\sigma_{\Sigma_{0}}^{\Sigma})_{!}$ as an object of ${\mathcal{P}M}^{S}_{K}$. Now define ${\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}$ and $M_{f,\Sigma_{0}}^{\Sigma}$ exactly as before, i.e., as the intersection of the kernels of elements of $I_{f}$. Note that if $\Sigma_{0}\subset\Sigma$, then $M_{f,\Sigma_{0}}^{\Sigma}=M_{f}^{\Sigma}$.
@@ 7.1. Galois cohomology
Below we shall define an $\mathcal{O}$}-submodule $H^{1}_{w}(G_{p},A)\subset H^{1}(G_{p},A)$ for each prime pp. For each finite set of primes $\Sigma\subset\Sigma_{0}$, we define
@@
We define $H^{1}_{w}(G_{\ell},A_{n})$ to be the set of elements of $H^{1}(G_{\ell},A_{n})$ corresponding to extensions in $\mathcal{F}\mathcal{R}$. Using the stability of $\mathcal{F}\mathcal{R}$ under direct sums, subobjects and quotients, one checks that $H^{1}_{w}(G_{\ell},A_{n})$ is an $\mathcal{O}$}-submodule, and furthermore that it is the preimage of $H^{1}_{w}(G_{\ell},A_{n+1})$ under the natural map. Now define $H^{1}_{w}(G_{p},A)=\displaystyle{\lim_{\longrightarrow}}\,H^{1}_{w}(G_{p},A_{n})$ for every prime pp.
@@
is then the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{\bf f}(G_{p},A)$ for all primes pp. If $p\neq\ell$, it is clear that
@@
Now we specialize to the case where $\rho_{1}=\rho_{2}$, each having dimension 22 over $K$ with $\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}D_{\mbox{\scriptsize crys}}(V_{i})$ =1=1 (so also $\ell>2$). We assume also that $L_{1}=L_{2}$, and let
@@
where ${\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}L_{1}$ denotes the kernel of the trace map $\operatorname{{\mathop{\rm End}\nolimits}}_{\mathcal{O}}L\to{\mathcal{O}}$. We let $A^{0}_{n}=A^{0}[\lambda^{n+1}]$, identify $A^{0}$ with a submodule of AA, and define $H^{1}_{w}(G_{p},A^{0}_{n})$, $H^{1}_{w}(G_{p},A^{0})$ and $H^{1}_{w}(G_{\mathbb{Q}},A^{0})$ as the preimage of the corresponding group for AA. From the proposition above, we get
@@ 7.2. An axiomatic formulation
Recall that if AA is an object of $\mathcal{C}$} with maximal ideal $\mathfrak{m}$}, then an AA-deformation of $V_{0}$ is an isomorphism class of free AA-modules $M$ endowed with continuous $AG_{\mathbb{Q}}$-action such that $M/{\mathfrak{m}}M$ is $(A/{\mathfrak{m}})G_{\mathbb{Q}}$-isomorphic to $(A/{\mathfrak{m}})\otimes_{\kappa}V_{0}$.
@@
Consider the functor on $\mathcal{C}$} which associates to AA the set of AA-deformations of $\rho_{0}$ of type $\Sigma$. By the results of Mazur and Ramakrishna, this functor is representable by an object of $\mathcal{C}$}. We denote this object $R^{\Sigma}$ and let $M^{\Sigma}$ denote the universal deformation. We recall also that $R^{\Sigma}$ is topologically generated by the elements $t_{g}^{\Sigma}$ for $g$ in $G_{\mathbb{Q}}$, where $t_{g}^{\Sigma}$ denotes the trace of the endomorphism $g$ of the free $R^{\Sigma}$-module $M^{\Sigma}$. In particular, $R^{\Sigma}$ has residue field $\kappa$.
@@ 7.3. The construction
Now suppose $\delta=1$. Let ${\mathcal{V}}_{0}={\mathcal{V}}_{0}^{\Sigma}={\mathcal{V}}_{0}^{\Sigma^{\prime}}$, $V^{\prime\prime}=\alpha_{p}^{-1}V^{\prime}\alpha_{p}$ and $A={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime\prime})$. Then $V^{\prime\prime}$ is normal in both $U^{\Sigma}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$, and both groups act on AA, where the action of $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ on ${\mathcal{V}}_{0}$ is defined by first conjugating by $\alpha_{p}$. Note that the actions agree on
@@
is injective. To prove this, suppose there were a nonzero vector in AA fixed by both $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ and $A^{U^{\Sigma}}$. Then there would be a nonzero vector in ${\mathcal{M}}_{!}(V^{\prime\prime})$ on which $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ acts via $\psi_{p}^{-1}(d)$ for all matrices in $U_{p}^{\mathop{\rm min}\nolimits}$ or $\alpha_{p}^{-1}U_{p}^{\mathop{\rm min}\nolimits}\alpha_{p}$. This contradicts the fact that $\psi_{p}$ has conductor $p^{c}$. It follows that the map
@@
Now let $A_{c}$, $B_{c}$, $C_{c}$ and $D_{c}$ denote the $\tilde{{\mathbb{T}}}-$modules defined in the same way as AA, $B$, $C$ and $D$, but using ${\mathcal{M}}_{c}$ instead of $\mathcal{M}$$_{!}$. Then the natural maps $C_{c}\to A_{c}^{V}$, $D_{c}\to B_{c}^{\tilde{V}}$ and $D_{c}\to A_{c}^{\tilde{V}}$ are isomorphisms. Since $\mathfrak{m}$} is not Eisenstein, it follows that so are the maps $C_{\mathfrak{m}}^{U^{\Sigma}}\to A_{\mathfrak{m}}^{U}$, $D_{\mathfrak{m}}^{\tilde{U}}\to B_{\mathfrak{m}}^{\tilde{U}}$ and $D_{\mathfrak{m}}^{\tilde{U}}\to A_{\mathfrak{m}}^{\tilde{U}}$. We now obtain the desired injectivity on combining the localization at $\mathfrak{m}$} of the exact sequences (47) and (48) (cf. the discussion after Lemma 2.5 of [Wi]). ∎
@@ 7.4. Consequences
Suppose that $K_{0}$ is a number field with ring of integers ${\mathcal{O}}_{0}$, and $f=\sum a_{n}e^{2\pi inz}$ is a newform of weight $k\geq 2$ and level NN with coefficients $a_{n}$ in $K_{0}$. For primes $\lambda$ of $K_{0}$, we write $\bar{\rho}_{f,\lambda}$ for the representation of $G_{\mathbb{Q}}$ on
@@
Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K_{0}$. Suppose that $\lambda$ is a prime in $S_{f}$ (as defined above), and let ${\mathcal{O}}={\mathcal{O}}_{0,\lambda}$ and $K=K_{0,\lambda}$. Suppose that $\Sigma$ is a finite set of primes not containing $\ell$ such that $M_{f,\lambda}$ is minimally ramified outside $\Sigma$. Then the $\mathcal{O}$}-module
@@ 8.2. Deligne’s period
Recall that $\Sigma_{e}(f)$ is the set of primes $p$ such that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=1$ but $L_{p}(A_{f},s)=(1+p^{-s})^{-1}$. We write $b_{{\mathop{\rm dR}\nolimits}}$ for the basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}$ defined in §1.3, and set $\eta=0$ or 11 so $\eta\equiv k\bmod 2$. Note that by Proposition 5.5 of [De3], we have $\epsilon(M_{f}\otimes M_{\psi^{-1}})/\epsilon(M_{\psi^{-1}})\in K^{\times}$. The same proposition together with Lemma 6.1 gives $\epsilon(A_{f})\in K^{\times}$.
@@
As in §4.5 put $U=U_{0}(N)$, let $\sigma:U\rightarrow K^{\times}$ be the representation $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\psi^{-1}(a_{N})$, and choose a multiple $N^{\prime}$ of NN so that $U_{N^{\prime}}\subseteq U$ is sufficiently small. Put $w=\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)\in GL_{2}({\mathbb{A}}_{\bf f})$ and denote by $W=[UwU]_{\omega}:M(\sigma)_{N^{\prime}}\rightarrow M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ the isomorphism in Lemma 5.2. Note that $ww_{N}^{-1}\in U$ so we can work with $w$ instead of $w_{N}$. For any one-dimensional $K$-representation $\sigma$ of $U$ we shall view $M(\sigma)_{N^{\prime}}$ as a sub-PM-structure of $K\otimes M_{N^{\prime}}$. With $I^{\infty}$ denoting the comparison isomorphism for both $M_{f}$ and $K\otimes M_{N^{\prime}}$ we have
@@
We proceed with the computation of $Wf\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Note that the field $K_{f}$ generated by the Fourier coefficients of the newform $f$ is either totally real or a CM field and hence has a well defined automorphism $\rho$ induced by complex conjugation. It is known that the Fourier expansion $f^{\rho}(z)=\sum_{n=1}^{\infty}a_{n}^{\rho}e^{2\pi izn}$ is a newform of conductor NN and character $\psi^{-1}$ [Miy, 4.6.15(2)], hence represents an element of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma})_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Recall that $b_{\mathop{\rm dR}\nolimits}$ is the canonical basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}\cong(H^{0}(X_{N^{\prime}})\otimes K_{\psi^{-1}})^{U}_{\mathop{\rm dR}\nolimits}$ (Lemma 4.6) and that $M_{\psi}\otimes_{K}M_{f^{\rho}}$ has a natural map into $M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ via the cup product on $X_{N^{\prime}}$ (see (26)).
@@
where $\phi$ is Euler’s function and we have used Lemma $2.$ Therefore (58) equals
@@
If $f$ is a newform of conductor NN, weight $k$ and with coefficients in the number field $K$ we have
@@
for all primes pp. If $p\nmid N$ this is immediate from §$5.$ If $N_{p}=p$ and $N_{\psi,p}=1$ we have $a_{p}^{2}=\psi(p)p^{k-2}$ by [Miy, Th. 4.6.17(2)] and $\pi_{p}(f)$ is special so that (61) holds true by (38). The only other case in which $a_{p}\neq 0$ is when $N_{p}=N_{\psi,p}$ [Miy, Th. 4.6.17]. In this case $\bar{\psi}(p)=0$ and hence $L_{p}(k,f,\bar{\psi})=1$ whereas $\pi_{p}(f)$ is principal series so that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=(1-p^{-1})^{-1}=N_{p}/\phi(N_{p})$ by (38). Finally, if $N_{p}>1$ and $a_{p}=0$ then $L_{p}(k,f,\bar{\psi})=L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=1$, $N_{p}/N_{\psi,p}>1$ and both sides in (61) equal $(1-p^{-1})$. ∎
@@
where $\epsilon_{0}$ is introduced in [De2, §5] and $(M_{\lambda}|W_{p})^{ss}$ is the semisimplification of $M_{\lambda}$ as a representation of $W_{p}$. Now for any $\lambda\nmid p$ the $W_{p}$-representation $M_{\lambda}$ is the restriction of a continuous $G_{p}$-representation, hence carries a $W_{p}$-stable ${\mathcal{O}}_{\lambda}$-lattice. This implies, on the one hand, that ${\mathop{\rm det}\nolimits}(-\operatorname{{\mathop{\rm Frob}\nolimits}}|M_{\lambda}^{I_{p}})\in{\mathcal{O}}_{\lambda}^{\times}$ and on the other hand, via [De2, Th. 6.5(c)], that $\epsilon_{0}((M_{\lambda}|W_{p})^{ss},\psi_{p},dx_{p})\in{\mathcal{O}}_{\lambda}[\mu_{p^{\infty}}]^{\times}$. Noting that with our choice of $\psi_{p},dx_{p}$ the epsilon factor equals 1 (resp. a power of ii) for $p\nmid c(M)$ (resp. $p=\infty$) the Lemma follows. ∎

### v2

@@ 6.5. Exceptional primes
where $V_{p}^{\prime}$ denotes the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to $I$ mod $p^{c_{p}/2}$, and $\Theta(\varepsilon)$ is defined in §3.2 of [C-D-T]. Our hypotheses ensure that $\varepsilon/\varepsilon\circ\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ has conductor $(p{\mathcal{O}}_{F})^{c_{p}/2}$. Let $g_{p}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{c_{p}/2}}}\end{array}\right)$}}$. We recall the following properties of $\Theta(\varepsilon)$:
@@
where ${\mathcal{V}}^{\Sigma}_{p,\Sigma_{0}}={\mathcal{V}}_{p}^{0}$ if $p\in\Sigma_{0}-\Sigma$ and ${\mathcal{V}}_{p}^{1}$ (restricted to $U_{p}$) otherwise. Here the tensor product is over all primes $p$, but the action is trivial for all but finitely many $p$. Note that if $\Sigma_{0}\subset\Sigma$, then $U^{\Sigma}_{\Sigma_{0}}=U^{\Sigma}$ and $\sigma^{\Sigma}_{\Sigma_{0}}=\sigma(N,\psi)$.
@@
It follows from the lemma and the definitions that $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$ is injective and commutes with the action of the Hecke operators $T_{p}$ for all primes $p$. In particular, for $p\in S$, the operator $T_{p}$ can be written as a polynomial in $T_{q}$ for $q\neq p$, and the same arguments as in Proposition 5.6 show that $\mathbb{T}$} acts on ${\mathcal{M}}(\sigma_{\Sigma_{0}}^{\Sigma})_{!}$ as an object of ${\mathcal{P}M}^{S}_{K}$. Now define ${\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}$ and $M_{f,\Sigma_{0}}^{\Sigma}$ exactly as before, i.e., as the intersection of the kernels of elements of $I_{f}$. Note that if $\Sigma_{0}\subset\Sigma$, then $M_{f,\Sigma_{0}}^{\Sigma}=M_{f}^{\Sigma}$.
@@ 7.1. Galois cohomology
Below we shall define an $\mathcal{O}$}-submodule $H^{1}_{w}(G_{p},A)\subset H^{1}(G_{p},A)$ for each prime $p$. For each finite set of primes $\Sigma\subset\Sigma_{0}$, we define
@@
We define $H^{1}_{w}(G_{\ell},A_{n})$ to be the set of elements of $H^{1}(G_{\ell},A_{n})$ corresponding to extensions in $\mathcal{F}\mathcal{R}$. Using the stability of $\mathcal{F}\mathcal{R}$ under direct sums, subobjects and quotients, one checks that $H^{1}_{w}(G_{\ell},A_{n})$ is an $\mathcal{O}$}-submodule, and furthermore that it is the preimage of $H^{1}_{w}(G_{\ell},A_{n+1})$ under the natural map. Now define $H^{1}_{w}(G_{p},A)=\displaystyle{\lim_{\longrightarrow}}\,H^{1}_{w}(G_{p},A_{n})$ for every prime $p$.
@@
is then the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{\bf f}(G_{p},A)$ for all primes $p$. If $p\neq\ell$, it is clear that
@@
Now we specialize to the case where $\rho_{1}=\rho_{2}$, each having dimension $2$ over $K$ with $\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}D_{\mbox{\scriptsize crys}}(V_{i})$ =1=1 (so also $\ell>2$). We assume also that $L_{1}=L_{2}$, and let
@@
where ${\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}L_{1}$ denotes the kernel of the trace map $\operatorname{{\mathop{\rm End}\nolimits}}_{\mathcal{O}}L\to{\mathcal{O}}$. We let $A^{0}_{n}=A^{0}[\lambda^{n+1}]$, identify $A^{0}$ with a submodule of $A$, and define $H^{1}_{w}(G_{p},A^{0}_{n})$, $H^{1}_{w}(G_{p},A^{0})$ and $H^{1}_{w}(G_{\mathbb{Q}},A^{0})$ as the preimage of the corresponding group for $A$. From the proposition above, we get
@@ 7.2. An axiomatic formulation
Recall that if $A$ is an object of $\mathcal{C}$} with maximal ideal $\mathfrak{m}$}, then an $A$-deformation of $V_{0}$ is an isomorphism class of free $A$-modules $M$ endowed with continuous $AG_{\mathbb{Q}}$-action such that $M/{\mathfrak{m}}M$ is $(A/{\mathfrak{m}})G_{\mathbb{Q}}$-isomorphic to $(A/{\mathfrak{m}})\otimes_{\kappa}V_{0}$.
@@
Consider the functor on $\mathcal{C}$} which associates to $A$ the set of $A$-deformations of $\rho_{0}$ of type $\Sigma$. By the results of Mazur and Ramakrishna, this functor is representable by an object of $\mathcal{C}$}. We denote this object $R^{\Sigma}$ and let $M^{\Sigma}$ denote the universal deformation. We recall also that $R^{\Sigma}$ is topologically generated by the elements $t_{g}^{\Sigma}$ for $g$ in $G_{\mathbb{Q}}$, where $t_{g}^{\Sigma}$ denotes the trace of the endomorphism $g$ of the free $R^{\Sigma}$-module $M^{\Sigma}$. In particular, $R^{\Sigma}$ has residue field $\kappa$.
@@ 7.3. The construction
Now suppose $\delta=1$. Let ${\mathcal{V}}_{0}={\mathcal{V}}_{0}^{\Sigma}={\mathcal{V}}_{0}^{\Sigma^{\prime}}$, $V^{\prime\prime}=\alpha_{p}^{-1}V^{\prime}\alpha_{p}$ and $A={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime\prime})$. Then $V^{\prime\prime}$ is normal in both $U^{\Sigma}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$, and both groups act on $A$, where the action of $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ on ${\mathcal{V}}_{0}$ is defined by first conjugating by $\alpha_{p}$. Note that the actions agree on
@@
is injective. To prove this, suppose there were a nonzero vector in $A$ fixed by both $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ and $A^{U^{\Sigma}}$. Then there would be a nonzero vector in ${\mathcal{M}}_{!}(V^{\prime\prime})$ on which $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ acts via $\psi_{p}^{-1}(d)$ for all matrices in $U_{p}^{\mathop{\rm min}\nolimits}$ or $\alpha_{p}^{-1}U_{p}^{\mathop{\rm min}\nolimits}\alpha_{p}$. This contradicts the fact that $\psi_{p}$ has conductor $p^{c}$. It follows that the map
@@
Now let $A_{c}$, $B_{c}$, $C_{c}$ and $D_{c}$ denote the $\tilde{{\mathbb{T}}}-$modules defined in the same way as $A$, $B$, $C$ and $D$, but using ${\mathcal{M}}_{c}$ instead of $\mathcal{M}$$_{!}$. Then the natural maps $C_{c}\to A_{c}^{V}$, $D_{c}\to B_{c}^{\tilde{V}}$ and $D_{c}\to A_{c}^{\tilde{V}}$ are isomorphisms. Since $\mathfrak{m}$} is not Eisenstein, it follows that so are the maps $C_{\mathfrak{m}}^{U^{\Sigma}}\to A_{\mathfrak{m}}^{U}$, $D_{\mathfrak{m}}^{\tilde{U}}\to B_{\mathfrak{m}}^{\tilde{U}}$ and $D_{\mathfrak{m}}^{\tilde{U}}\to A_{\mathfrak{m}}^{\tilde{U}}$. We now obtain the desired injectivity on combining the localization at $\mathfrak{m}$} of the exact sequences (47) and (48) (cf. the discussion after Lemma 2.5 of [Wi]). ∎
@@ 7.4. Consequences
Suppose that $K_{0}$ is a number field with ring of integers ${\mathcal{O}}_{0}$, and $f=\sum a_{n}e^{2\pi inz}$ is a newform of weight $k\geq 2$ and level $N$ with coefficients $a_{n}$ in $K_{0}$. For primes $\lambda$ of $K_{0}$, we write $\bar{\rho}_{f,\lambda}$ for the representation of $G_{\mathbb{Q}}$ on
@@
Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K_{0}$. Suppose that $\lambda$ is a prime in $S_{f}$ (as defined above), and let ${\mathcal{O}}={\mathcal{O}}_{0,\lambda}$ and $K=K_{0,\lambda}$. Suppose that $\Sigma$ is a finite set of primes not containing $\ell$ such that $M_{f,\lambda}$ is minimally ramified outside $\Sigma$. Then the $\mathcal{O}$}-module
@@ 8.2. Deligne’s period
Recall that $\Sigma_{e}(f)$ is the set of primes $p$ such that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=1$ but $L_{p}(A_{f},s)=(1+p^{-s})^{-1}$. We write $b_{{\mathop{\rm dR}\nolimits}}$ for the basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}$ defined in §1.3, and set $\eta=0$ or $1$ so $\eta\equiv k\bmod 2$. Note that by Proposition 5.5 of [De3], we have $\epsilon(M_{f}\otimes M_{\psi^{-1}})/\epsilon(M_{\psi^{-1}})\in K^{\times}$. The same proposition together with Lemma 6.1 gives $\epsilon(A_{f})\in K^{\times}$.
@@
As in §4.5 put $U=U_{0}(N)$, let $\sigma:U\rightarrow K^{\times}$ be the representation $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\psi^{-1}(a_{N})$, and choose a multiple $N^{\prime}$ of $N$ so that $U_{N^{\prime}}\subseteq U$ is sufficiently small. Put $w=\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)\in GL_{2}({\mathbb{A}}_{\bf f})$ and denote by $W=[UwU]_{\omega}:M(\sigma)_{N^{\prime}}\rightarrow M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ the isomorphism in Lemma 5.2. Note that $ww_{N}^{-1}\in U$ so we can work with $w$ instead of $w_{N}$. For any one-dimensional $K$-representation $\sigma$ of $U$ we shall view $M(\sigma)_{N^{\prime}}$ as a sub-PM-structure of $K\otimes M_{N^{\prime}}$. With $I^{\infty}$ denoting the comparison isomorphism for both $M_{f}$ and $K\otimes M_{N^{\prime}}$ we have
@@
We proceed with the computation of $Wf\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Note that the field $K_{f}$ generated by the Fourier coefficients of the newform $f$ is either totally real or a CM field and hence has a well defined automorphism $\rho$ induced by complex conjugation. It is known that the Fourier expansion $f^{\rho}(z)=\sum_{n=1}^{\infty}a_{n}^{\rho}e^{2\pi izn}$ is a newform of conductor $N$ and character $\psi^{-1}$ [Miy, 4.6.15(2)], hence represents an element of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma})_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Recall that $b_{\mathop{\rm dR}\nolimits}$ is the canonical basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}\cong(H^{0}(X_{N^{\prime}})\otimes K_{\psi^{-1}})^{U}_{\mathop{\rm dR}\nolimits}$ (Lemma 4.6) and that $M_{\psi}\otimes_{K}M_{f^{\rho}}$ has a natural map into $M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ via the cup product on $X_{N^{\prime}}$ (see (26)).
@@
where $\phi$ is Euler’s function and we have used Lemma 2.2. Therefore (58) equals
@@
If $f$ is a newform of conductor $N$, weight $k$ and with coefficients in the number field $K$ we have
@@
for all primes $p$. If $p\nmid N$ this is immediate from §5.5. If $N_{p}=p$ and $N_{\psi,p}=1$ we have $a_{p}^{2}=\psi(p)p^{k-2}$ by [Miy, Th. 4.6.17(2)] and $\pi_{p}(f)$ is special so that (61) holds true by (38). The only other case in which $a_{p}\neq 0$ is when $N_{p}=N_{\psi,p}$ [Miy, Th. 4.6.17]. In this case $\bar{\psi}(p)=0$ and hence $L_{p}(k,f,\bar{\psi})=1$ whereas $\pi_{p}(f)$ is principal series so that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=(1-p^{-1})^{-1}=N_{p}/\phi(N_{p})$ by (38). Finally, if $N_{p}>1$ and $a_{p}=0$ then $L_{p}(k,f,\bar{\psi})=L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=1$, $N_{p}/N_{\psi,p}>1$ and both sides in (61) equal $(1-p^{-1})$. ∎
@@
where $\epsilon_{0}$ is introduced in [De2, §5] and $(M_{\lambda}|W_{p})^{ss}$ is the semisimplification of $M_{\lambda}$ as a representation of $W_{p}$. Now for any $\lambda\nmid p$ the $W_{p}$-representation $M_{\lambda}$ is the restriction of a continuous $G_{p}$-representation, hence carries a $W_{p}$-stable ${\mathcal{O}}_{\lambda}$-lattice. This implies, on the one hand, that ${\mathop{\rm det}\nolimits}(-\operatorname{{\mathop{\rm Frob}\nolimits}}|M_{\lambda}^{I_{p}})\in{\mathcal{O}}_{\lambda}^{\times}$ and on the other hand, via [De2, Th. 6.5(c)], that $\epsilon_{0}((M_{\lambda}|W_{p})^{ss},\psi_{p},dx_{p})\in{\mathcal{O}}_{\lambda}[\mu_{p^{\infty}}]^{\times}$. Noting that with our choice of $\psi_{p},dx_{p}$ the epsilon factor equals 1 (resp. a power of $i$) for $p\nmid c(M)$ (resp. $p=\infty$) the Lemma follows. ∎